OpenStudy (erinweeks):

Write an equation in point-slope form of the line through point J(4, –4) with slope 4.

OpenStudy (anonymous):

This isn't Geometry, it's Pre-Algebra.

OpenStudy (erinweeks):

A. y+4=4(x+4) B. y+4=4(x-4) C. y-4=4(x+4) D. y+4=-4(x-4)

OpenStudy (erinweeks):

well im in geometry and this is a question on my homework!

OpenStudy (anonymous):

the answer is y + 4 =4(x+4), B.

OpenStudy (erinweeks):

thank you (:

OpenStudy (anonymous):

No problem.

OpenStudy (erinweeks):

hold on B ? you did a plus at (x+4) that would be A ?

OpenStudy (anonymous):

my bad I said the wrong thing, sorry

OpenStudy (anonymous):

you can take away my medal if you want.